OpenStudy (anonymous):

q+ 4 q ------- + --------- = 2 q−1 q+1 show all work please

OpenStudy (johnweldon1993):

First we need a common denominator to add these fractions...what would that be?

OpenStudy (anonymous):

would it be (q-1)(q+10) ??

OpenStudy (johnweldon1993):

Can I assume that 10 was a typo for 1?

OpenStudy (anonymous):

oops yeah it was

OpenStudy (johnweldon1993):

No problem, just want to make sure you got it :) \[\large \frac{(q+4)(q+1)}{(q-1)(q+1)} + \frac{q(q-1)}{(q-1)(q+1)} = 2\] Do you see how I got that with the common denominator?

OpenStudy (anonymous):

yup! so now would it be q^2+q+4q+4+q^2-q = 2q+2 +2q - 2 right ?

OpenStudy (johnweldon1993):

Not quite! \[\large \frac{q^2 + 5q + 4 + q^2 - q}{(q-1)(q+1)} = 2\] \[\large q^2 + 5q + 4 + q^2 - q = 2\color\red{(q-1)(q+1)}\] \[\large (q-1)(q+1) = q^2 - 1\] so \[\large 2q^2 + 4q + 4 = 2q^2 - 2\]